# Diffusion Models in Simulation-Based Inference: A Tutorial Review

### v1

@@ 3 Diffusion Models for Simulation-Based Inference
which approximates the trace of the Jacobian of the velocity field. The ODE formulation is particularly useful for SBI applications that require likelihood or evidence estimation, such as Bayesian model comparison [gunes2025compass].
Figure 4: Inverse kinematics toy example. Each panel shows 1,000 approximate posterior samples representing possible arm configurations $\boldsymbol{\theta}$ given an end-effector position $\mathbf{y}_{\text{obs}}$ (indicated by the red crosshair). The first panel shows the reference distribution estimated with ABC-SMC in pyABC [schaelte2022pyabc], while the remaining panels display samples from nine different diffusion models (see Section 5 for more details). The white segments indicate maximum a posteriori estimates. The quality of the posterior samples differs widely between the models, as also captured by differences in the maximum mean discrepancy (MMD) metric computed against the ABC posterior.
@@
In the previous sections, we introduced diffusion models for conditional density estimation in SBI. We now turn to a concrete application and demonstrate how these concepts are used in practice. We set the stage with the inverse kinematics problem from [kruse2021benchmarking], which provides an intuitive and visually accessible test case for SBI (Figure 4). The goal is to reconstruct the configuration of a planar robot arm from the observed position of its end effector. The unknown configuration is described by a four-dimensional parameter vector $\boldsymbol{\theta}\in\mathbb{R}^{4}$, where $\theta_{1}$ denotes the vertical offset of the arm’s base and $\theta_{2}$, $\theta_{3}$, $\theta_{4}$ are the angles at the three joints. Given a configuration $\boldsymbol{\theta}$, a deterministic simulator returns the two-dimensional end-effector position $\mathbf{y}_{\text{obs}}\in\mathbb{R}^{2}$.
@@
We train diffusion models with different design choices, discussed in detail in Section 5, such as the noise schedule and the learning target (e.g., the score $s$ or the vector field $v$), using simulated pairs $(\boldsymbol{\theta},\mathbf{y})$. Training proceeds by gradually diffusing samples of $\boldsymbol{\theta}$ and learning to reverse this process while conditioning on the observation $\mathbf{y}$. At inference time, we obtain posterior draws by running the reverse SDE or probability-flow ODE from an initial noise draw toward $\boldsymbol{\theta}$, guided by the position $\mathbf{y}_{\text{obs}}=(0,1.5)$. In our experiments, we train nine diffusion-based posterior estimators under a limited simulation budget of B=10, [^0] B=10{,}000, keeping network architecture and training hyperparameters fixed across models (Section A.1). For reference, we additionally approximate a high-fidelity posterior using ABC-SMC implemented in pyABC [schaelte2022pyabc] with a minimum acceptance threshold of $\epsilon_{\text{ABC}}=0.002$.
Under the same restricted simulation budget, the diffusion models exhibit noticeably different posterior quality, both visually and in terms of maximum mean discrepancy [gretton2012kernel, MMD;] (Figure 4). All models recover the characteristic multi-modality of the inverse kinematics problem by producing distinct, plausible arm configurations that reach the same end-effector position. Yet, differences between the posteriors are clearly visible when visualizing posterior samples as arm configurations and are reflected quantitatively in higher MMD values (Figure 4). The observed variability highlights the need for a systematic analysis of diffusion model design decisions for SBI, which we introduce in Section 5 and evaluate in Section 6. Before turning to these design considerations, however, we first examine how learned score functions can be exploited beyond standard posterior or likelihood estimation, since it is precisely this flexibility at training and inference time that makes diffusion models particularly powerful for SBI.
@@ 4 Special Use Cases of the Score: Compositional and Guided Inference
NPE & NLE [simons2023neural] high-quality samples VE linear, SDE, score
@@
[sharrock2024sequential] sequential correction in score space, high quality for single observation VP/VE linear, ODE, score
[Yu2025-neuralFidelityCalibrationInformativeSim-to-realAdaptationf] multiple prediction tasks with one model VP/VE linear, ODE, score
@@
[liang2025recent] high quality samples & high dimensional flow matching / consistency model
NPE [geffner2023compositional] compositional score for complete pooling VP linear, Langevin, score
[wildberger2023flow] high-dimensional & complex posteriors flow matching (power-law)
@@
[Andry2023-dataAssimilationSimulation-basedInferences] decomposing score for time varying prior/posterior & state space models VP, SDE, $\boldsymbol{\epsilon}$-prediction
[Pidstrigach2023-infinite-dimensionalDiffusionModelsq] infinite-dimensional score for non-parametric inference VP linear, SDE, score
@@
[linhart2024diffusion] compositional score for complete pooling VP linear, $\boldsymbol{\epsilon}$-prediction
@@
[holzschuh2024improving] decomposing flow for high-quality samples with simulator feedback flow matching
[orozco2024machine] high-quality samples VE-EDM, SDE, $\mathbf{F}$-prediction
[Fluri2024-improvingFlowMatchingSimulation-basedInferenceu] high-quality samples OT flow matching
[schmitt2024consistency] single-step sampling VE EDM, consistency model
[dasgupta2025conditional] high-dimensional VP linear, Langevin, score
[gebhard2025flow] high-quality samples flow matching ($\sqrt[4]{t}\sim\mathcal{U}[0,1]$)
[gloeckler2025compositional] compositional score for time series pooling VP linear, SDE, score
[chen2025conditional] high-quality samples VE EDM, ODE, $\mathbf{F}$-prediction
[yang2025priorguide] score guidance for post-hoc prior change VE EDM, SDE, score
[Zeng2025-full-waveformVariationalInferenceFullCommon-imageGathersDiffusionNetworkn] high-dimensional EDM, SDE, $\mathbf{F}$-prediction
[Zeng2025-amortizedLikelihood-freeStochasticModelUpdatingViaConditionalDiffusionModelStructuralHealthMonitoringw] high-quality samples sub-VP linear, ODE, score
@@
[Nautiyal2025-condisimConditionalDiffusionModelsSimulationBasedInferencef] high-quality samples VP cosine/quadratic/linear, $\boldsymbol{\epsilon}$-prediction
[dirmeier2025causal] causal network architecture flow matching
[arruda2025compositional] compositional score for hierarchical models VP cosine, SDE, $\boldsymbol{v}$-prediction
@@
[cirakman2025efficient] multi-scale wavelet score composition for high-dimensional problems VE EDM, SDE, $\mathbf{F}$-prediction
[Raymond2025-sbiIMBHB] high-quality samples flow matching
[orsini2025flow] high-quality samples flow matching (power-law)
[zhao2025generativeconsistencymodelsestimation] fast & high-quality samples VE EDM, consistency model
[orsini2025flow2] high-quality samples flow matching (power-law)
@@
[touron2025erroranalysiscompositionalscorebased] compositional score for complete pooling VP linear, $\boldsymbol{\epsilon}$-prediction
[ruhlmann2025flow] robust inference under model misspecification flow matching
[viterbo2025dynamical] high-quality samples flow matching
[ko2025latent] simulation efficiency using tractable score VP linear, SDE, score
@@
[villarreal2025machine] high-quality samples flow matching & VE linear, SDE, $\boldsymbol{\epsilon}$-prediction
@@
NLE [haitsiukevich2024diffusion] multiple prediction tasks with PDEs VP linear/VE EDM, ODE, $\boldsymbol{\epsilon}$- & $\mathbf{F}$-prediction
[shysheya2024conditional] forecasting PDE VP linear, SDE, score
[Mudur2025-diffusion-hmcParameterInferenceDiffusion-model-drivenHamiltonianMonteCarlob] high-dimensional VP linear, SDE, $\boldsymbol{\epsilon}$-prediction
[Legin2025-gravitational-waveParameterEstimationNon-gaussianNoiseUsingScore-basedLikelihoodCharacterizationl] learn complex noise model VE linear, score
@@
Joint [gloeckler2024all] score decomposition to sample arbitrary conditionals & guidance for post-hoc adaptation to new priors/likelihoods VP/VE linear, SDE, score
[Tesso2025-spatformerSimulation-basedInferenceTransformersSpatialStatisticsi] non-parametric spatial inference VE linear, SDE, score
@@
[gunes2025compass] model comparison using likelihood and inference with posterior VE linear, SDE/ODE, score
[jeong2025conditionalflowmatchingbayesian] theoretical guarantee on consistency flow matching
[charles2025tokenised] hierarchical models flow matching
@@ 4.1.1 Guiding the Reverse Process
In conditional image generation, a widely used method known as classifier guidance [dhariwal2021diffusion] approximates the likelihood score via an auxiliary classifier that predicts class labels from noisy samples. Its gradient with respect to the input is then added to the reverse process, and the resulting samples are steered toward the target class-conditional distribution. Classifier guidance is rarely useful for SBI, where the conditions are typically continuous quantities rather than discrete labels.
A more flexible method is classifier-free guidance [Ho2022-classifier-freeDiffusionGuidancez]. Instead of relying on an external classifier, the diffusion model is trained jointly in two randomly switching modes: conditional (with access to the conditioning variables, i.e., the posterior) and unconditional (with the conditioning masked out, i.e., the prior). The likelihood score can be estimated from the difference between the conditional and unconditional scores as
@@
where the scales ($\alpha_{1},\alpha_{2}$) and optional shifts ($\beta_{1},\beta_{2}$) act as guidance dials, allowing tempering and shifting of the prior and likelihood contributions [gloeckler2024all]. In theory, these dials enable tempered-like posterior sampling and incorporation of side information, since the user can control how strongly the model adheres to the conditions. Such modifications can be used to perform sensitivity analysis with respect to prior and likelihood choices [elsemuller2023sensitivity], or to reweigh parts of the likelihood in order to emphasize specific subsets of the data, such as recent or high-quality observations.
Notably, guidance is not limited to labels or direct conditioning variables (Figure 5). Recent works have shown that diffusion models can be guided by almost arbitrary functions of the generated sample [bansal2023universal] or even surrogates of their scores in cases of non-differentiability [Shen2024-moleculesnondiffguidance]. This makes it possible to design and impose custom constraints during the reverse process without retraining the model. For instance, one can require that generated samples satisfy inequality constraints or lie within prescribed intervals. Consider imposing $K$ constraints of the form $c_{k}(\mathbf{z}_{t})\leq 0$. One can incorporate them directly into the score estimate by adding a correction term derived from the gradient of a soft penalty. A possible formulation is
@@
where $\operatorname{sigmoid}$ denotes the logistic sigmoid and $h(t)$ is a scaling function that diverges as $t\to 0$, typically linked to the noise schedule. This ensures that as the process approaches the target sample, the constraints are enforced with increasing strength. Similarly, the posterior can be conditioned on observation intervals or other context variables [gloeckler2024all].
However, the flexibility that makes guidance attractive also carries a cost: any guidance biases the reverse diffusion process [chidambaram2024does]. Changing the score means that one implicitly assumes that the marginal density $p(\mathbf{z}_{t}\mid\mathbf{z}_{0})$ of the reverse diffusion process changes accordingly. Depending on the type of guidance, this can make the new score unstable and require Langevin or MCMC-based samplers to correct for the error [geffner2023compositional, sjoberg2023mcmc]. However, the need for correction depends on the application and the guidance strength. Moreover, [vuong2025we] argue that diffusion models can be reinterpreted through Wasserstein gradient flow matching: terminal samples may follow the altered marginal, even if the stochastic process along the way does not replicate the true time reversal of diffusion.
@@ 4.1.2 Incorporating Simulator Feedback
So far, guidance signals have only been injected at inference time: the total score used in the reverse process is steered toward desirable properties without retraining the model. A complementary strategy is to integrate feedback during training. In the flow-matching setting of [holzschuh2024improving], this leads to a two-phase procedure.
In the first phase, a primary flow is trained to estimate a vector field $\hat{\mathbf{u}}$ that transports noise to parameters via the reverse probability flow (8). In addition to conditioning on $\mathbf{y}$, a form of self-conditioning [chen2023-selfconditioning] is introduced to improve the posterior approximation. At each time $t<0.5$, where we are closer to the clean target in the diffusion process, the model makes a one-step prediction of the clean state $\tilde{\boldsymbol{\theta}}_{0}$ from the current noisy state $\boldsymbol{\theta}_{t}$ and then self-conditions the velocity field on that prediction:
@@
For diffusion models, in cases where the score $\nabla_{\boldsymbol{\theta}_{t}}\!\log p(\boldsymbol{\theta}_{t}\mid\mathbf{y})$ is tractable, incorporating the score in the objective to provide an unbiased estimator of the marginal score leads to higher simulation efficiency [ko2025latent]. Beyond SBI, related work has also explored incorporating feedback into flow-based models by differentiating through the flow and directly optimizing a loss to find the optimal source noise [ben2024d], or by guiding diffusion models to satisfy a terminal cost [pandey2025variational]. In such approaches, the loss could be likelihood-based, providing an alternative form of simulator feedback that modifies the generative process via optimization rather than through additional observation-specific networks.

### v2

@@ 3 Diffusion Models for Simulation-Based Inference
which approximates the trace of the Jacobian of the velocity field. The ODE formulation is particularly useful for SBI applications that require likelihood or evidence estimation, such as Bayesian model comparison [63].
Figure 4: Inverse kinematics toy example. Each panel shows 1,000 approximate posterior samples representing possible arm configurations $\boldsymbol{\theta}$ given an end-effector position $\mathbf{y}_{\text{obs}}$ (indicated by the red crosshair). The first panel shows the reference distribution estimated with ABC-SMC in pyABC [149], while the remaining panels display samples from nine different diffusion models (see Section 5 for more details). The white segments indicate maximum a posteriori estimates. The quality of the posterior samples differs widely between the models, as also captured by differences in the maximum mean discrepancy (MMD) metric computed against the ABC posterior.
@@
In the previous sections, we introduced diffusion models for conditional density estimation in SBI. We now turn to a concrete application and demonstrate how these concepts are used in practice. We set the stage with the inverse kinematics problem from [97], which provides an intuitive and visually accessible test case for SBI (Figure 4). The goal is to reconstruct the configuration of a planar robot arm from the observed position of its end effector. The unknown configuration is described by a four-dimensional parameter vector $\boldsymbol{\theta}\in\mathbb{R}^{4}$, where $\theta_{1}$ denotes the vertical offset of the arm’s base and $\theta_{2}$, $\theta_{3}$, $\theta_{4}$ are the angles at the three joints. Given a configuration $\boldsymbol{\theta}$, a deterministic simulator returns the two-dimensional end-effector position $\mathbf{y}_{\text{obs}}\in\mathbb{R}^{2}$.
@@
We train diffusion models with different design choices, discussed in detail in Section 5, such as the noise schedule and the learning target (e.g., the score $s$ or the vector field $v$), using simulated pairs $(\boldsymbol{\theta},\mathbf{y})$. Training proceeds by gradually diffusing samples of $\boldsymbol{\theta}$ and learning to reverse this process conditioned on the observation $\mathbf{y}_{\text{obs}}$. At inference time, we obtain posterior draws by running the reverse SDE or probability-flow ODE from an initial noise draw toward $\boldsymbol{\theta}$, guided by the position $\mathbf{y}_{\text{obs}}=(0,1.5)$. For this illustration, we train nine diffusion-based posterior estimators under a limited simulation budget of B=10, [^0] B=10{,}000, keeping network architecture and training hyperparameters fixed across models (Section A.1). For reference, we additionally approximate a high-fidelity posterior using ABC-SMC implemented in pyABC [149] with a minimum acceptance threshold of $\epsilon_{\text{ABC}}=0.002$.
Under the same restricted simulation budget, the diffusion models exhibit noticeably different posterior quality, both visually and in terms of maximum mean discrepancy [61, MMD;] (Figure 4). All models recover the characteristic multi-modality of the inverse kinematics problem by producing distinct, plausible arm configurations that reach the same end-effector position. Yet, differences between the posteriors are clearly visible when visualizing posterior samples as arm configurations and are reflected quantitatively in higher MMD values (Figure 4). The observed variability highlights the need for a systematic analysis of diffusion model design decisions for SBI, which we introduce in Section 5 and evaluate in Section 6. Before turning to these design considerations, however, we first examine how learned score functions can be exploited beyond standard posterior or likelihood estimation, since it is precisely this flexibility at training and inference time that makes diffusion models particularly powerful for SBI.
@@ 4 Special Use Cases of the Score: Compositional and Guided Inference
NPE & NLE [156] high-quality samples VE linear, SDE, score
@@
[153] sequential correction in score space, high quality for single observation VP/VE linear, ODE, score
[195] multiple prediction tasks with one model VP/VE linear, ODE, score
@@
[104] high quality samples & high dimensional flow matching / consistency model
NPE [53] compositional score for complete pooling VP linear, Langevin, score
[190] high-dimensional & complex posteriors flow matching (power-law)
@@
[3] decomposing score for time varying prior/posterior & state space models VP, SDE, $\boldsymbol{\epsilon}$-prediction
[135] infinite-dimensional score for non-parametric inference VP linear, SDE, score
@@
[105] compositional score for complete pooling VP linear, $\boldsymbol{\epsilon}$-prediction
@@
[76] decomposing flow for high-quality samples with simulator feedback flow matching
[125] high-quality samples VE-EDM, SDE, $\mathbf{F}$-prediction
[47] high-quality samples OT flow matching
[151] single-step sampling VE EDM, consistency model
[34] high-dimensional VP linear, Langevin, score
[52] high-quality samples flow matching ($\sqrt[4]{t}\sim\mathcal{U}[0,1]$)
[59] compositional score for time series pooling VP linear, SDE, score
[23] high-quality samples VE EDM, ODE, $\mathbf{F}$-prediction
[192] score guidance for post-hoc prior change VE EDM, SDE, score
[199] high-dimensional EDM, SDE, $\mathbf{F}$-prediction
[198] high-quality samples sub-VP linear, ODE, score
@@
[121] high-quality samples VP cosine/quadratic/linear, $\boldsymbol{\epsilon}$-prediction
[40] causal network architecture flow matching
[4] compositional score for hierarchical models VP cosine, SDE, $\boldsymbol{v}$-prediction
@@
[29] multi-scale wavelet score composition for high-dimensional problems VE EDM, SDE, $\mathbf{F}$-prediction
[142] high-quality samples flow matching
[128] high-quality samples flow matching (power-law)
[204] fast & high-quality samples VE EDM, consistency model
[127] high-quality samples flow matching (power-law)
@@
[176] compositional score for complete pooling VP linear, $\boldsymbol{\epsilon}$-prediction
[144] robust inference under model misspecification flow matching
[182] high-quality samples flow matching
[96] simulation efficiency using tractable score VP linear, SDE, score
@@
[180] high-quality samples flow matching & VE linear, SDE, $\boldsymbol{\epsilon}$-prediction
@@
NLE [66] multiple prediction tasks with PDEs VP linear/VE EDM, ODE, $\boldsymbol{\epsilon}$- & $\mathbf{F}$-prediction
[155] forecasting PDE VP linear, SDE, score
[120] high-dimensional VP linear, SDE, $\boldsymbol{\epsilon}$-prediction
[101] learn complex noise model VE linear, score
@@
Joint [58] score decomposition to sample arbitrary conditionals & guidance for post-hoc adaptation to new priors/likelihoods VP/VE linear, SDE, score
[174] non-parametric spatial inference VE linear, SDE, score
@@
[63] model comparison using likelihood and inference with posterior VE linear, SDE/ODE, score
[85] theoretical guarantee on consistency flow matching
[20] hierarchical models flow matching
@@ 4.1.1 Guiding the Reverse Process
In conditional image generation, a widely used method known as classifier guidance [37] approximates the likelihood score via an auxiliary classifier that predicts class labels from noisy samples. Its gradient with respect to the input is then added to the reverse process, and the resulting samples are steered toward the target class-conditional distribution. Classifier guidance is rarely useful for SBI, where the conditions are typically continuous quantities rather than discrete labels.
A more flexible method is classifier-free guidance [73]. Instead of relying on an external classifier, the diffusion model is trained jointly in two randomly switching modes: conditional (with access to the conditioning variables, i.e., the posterior) and unconditional (with the conditioning masked out, i.e., the prior). The likelihood score can be estimated from the difference between the conditional and unconditional scores as
@@
where the scales ($\alpha_{1},\alpha_{2}$) and optional shifts ($\beta_{1},\beta_{2}$) act as guidance dials, allowing tempering and shifting of the prior and likelihood contributions [58]. In theory, these dials enable tempered-like posterior sampling and incorporation of side information, since the user can control how strongly the model adheres to the conditions. Such modifications can be used to perform sensitivity analysis with respect to prior and likelihood choices [42], or to reweigh parts of the likelihood in order to emphasize specific subsets of the data, such as recent or high-quality observations.
Notably, guidance is not limited to labels or direct conditioning variables (Figure 5). Recent works have shown that diffusion models can be guided by almost arbitrary functions of the generated sample [9] or even surrogates of their scores in cases of non-differentiability [154]. This makes it possible to design and impose custom constraints during the reverse process without retraining the model. For instance, one can require that generated samples satisfy inequality constraints or lie within prescribed intervals. Consider imposing $K$ constraints of the form $c_{k}(\mathbf{z}_{t})\leq 0$. One can incorporate them directly into the score estimate by adding a correction term derived from the gradient of a soft penalty. A possible formulation is
@@
where $\operatorname{sigmoid}$ denotes the logistic sigmoid and $h(t)$ is a scaling function that diverges as $t\to 0$, typically linked to the noise schedule, e.g., $h(t)=\alpha_{t}^{2}/\sigma_{t}^{2}$. This ensures that as the process approaches the target sample, the constraints are enforced with increasing strength. Similarly, the posterior can be conditioned on observation intervals or other context variables [58].
However, the flexibility that makes guidance attractive also carries a cost: any guidance biases the reverse diffusion process [27]. Changing the score means that one implicitly assumes that the marginal density $p(\mathbf{z}_{t}\mid\mathbf{z}_{0})$ of the reverse diffusion process changes accordingly. Depending on the type of guidance, this can make the new score unstable and require Langevin or MCMC-based samplers to correct for the error [53, 158]. However, the need for correction depends on the application and the guidance strength. Moreover, [184] argue that diffusion models can be reinterpreted through Wasserstein gradient flow matching: terminal samples may follow the altered marginal, even if the stochastic process along the way does not replicate the true time reversal of diffusion.
@@ 4.1.2 Incorporating Simulator Feedback
So far, guidance signals have only been injected at inference time: the total score used in the reverse process is steered toward desirable properties without retraining the model. A complementary strategy is to integrate feedback during training. In the flow-matching setting of [76], this leads to a two-phase procedure.
In the first phase, a primary flow is trained to estimate a vector field $\hat{\mathbf{u}}$ that transports noise to parameters via the reverse probability flow (8). In addition to conditioning on $\mathbf{y}$, a form of self-conditioning [24] is introduced to improve the posterior approximation. At each time $t<0.5$, where we are closer to the clean target in the diffusion process, the model makes a one-step prediction of the clean state $\tilde{\boldsymbol{\theta}}_{0}$ from the current noisy state $\boldsymbol{\theta}_{t}$ and then self-conditions the velocity field on that prediction:
@@
For diffusion models, in cases where the score $\nabla_{\boldsymbol{\theta}_{t}}\!\log p(\boldsymbol{\theta}_{t}\mid\mathbf{y})$ is tractable, incorporating the score in the objective to provide an unbiased estimator of the marginal score leads to higher simulation efficiency [96]. Beyond SBI, related work has also explored incorporating feedback into flow-based models by differentiating through the flow and directly optimizing a loss to find the optimal source noise [14], or by guiding diffusion models to satisfy a terminal cost [131]. In such approaches, the loss could be likelihood-based, providing an alternative form of simulator feedback that modifies the generative process via optimization rather than through additional observation-specific networks.